OpenStudy (anonymous):

Let G(x)=-1/3x+2 find 2g(7) Can somenoe please help me?

OpenStudy (anonymous):

can you find g(7) ?

OpenStudy (anonymous):

g(7)=-1/3?

OpenStudy (anonymous):

G(x) = -1/(3x) + 2 ?

OpenStudy (anonymous):

plug x = 7 into tihs equation and get g(7)

OpenStudy (anonymous):

 sorry

OpenStudy (anonymous):

ok so g(7) = -1/(3*7) + 2

OpenStudy (anonymous):

g(7)=-7/3 + 2?

OpenStudy (anonymous):

why ? g(7) = -1/(3*7) + 2 g(7) = -1/21 + 2

OpenStudy (anonymous):

the x is in the denominator

OpenStudy (anonymous):

okay so it is -1/21 + 2 so the answer is 1 20/21?

OpenStudy (anonymous):

for g(7) but you need 2g(7)